A container is lifted up 17 meters in 7 minutes. Another container of the same mass is lifted up twice as high in half the time. How much more power is needed to lift the second container as compared to the first one?
Answer #1

Power = Work done / time

= F*S / t

so P2/P1 = (S2/S1) * (t1 / t2)

= (2)*(2) = 4

so ratio of power needed to lift second to the power needed to lift the first one is 4